# Multi-News+: Cost-efficient Dataset Cleansing via LLM-based Data Annotation

### v1

@@ Abstract
In this study, we present a case study that extends the application of LLM-based data annotation to enhance the quality of existing datasets through a cleansing strategy. Specifically, we leverage approaches such as chain-of-thought (CoT) and majority voting to imitate human annotation and classify unrelated documents from the Multi-News dataset, which is widely used for the multi-document summarization task. Through our proposed cleansing method, we introduce an enhanced Multi-News+. By employing LLMs for data cleansing, we demonstrate an efficient and effective approach to improving dataset quality without relying on expensive human annotation efforts.
@@ 1 Introduction
The significance of dataset quality in deep learning applications cannot be overstated as mislabeled or noisy data can severely degrade performance Song et al. (2023). Datasets with incorrect labels, noise, or inconsistencies undermine the consistency and stability of model training. Cleansing these datasets contributes to enhancing model performance and generalization capabilities. Hence, ensuring the quality of the dataset by identifying and eliminating noisy data is imperative. In the realm of natural language processing, several researchers have attempted to improve the quality of noisy datasets. For example, ReDocRED Tan et al. (2022) addressed issues such as false negatives in DocRED Yao et al. (2019), a widely used dataset for relation extraction. Similarly, annotation inconsistencies were found in the MultiWOZ dataset Budzianowski et al. (2018) for dialogue state tracking Qian et al. (2021), leading to efforts to rectify these issues Eric et al. (2020); Zang et al. (2020); Han et al. (2021); Ye et al. (2022a).
@@ 3 Multi-News+
The previous Multi-News dataset plays an important role in multi-document summarization research. It consists of sets of documents and their corresponding summaries. However, as shown in Table 1 and detailed in Appendix F and G, the Multi-News dataset contains several noisy and irrelevant articles that are unrelated to the summary or other documents. This issue arises from their construction process, which relies on automated crawling from the Internet Archive.
To solve this issue and cleanse the dataset, we defined our problem as a classification task determining whether each document is relevant to the summary. To this end, we designed the prompt for the model as shown in Appendix I. We integrated CoT to enhance the model’s performance by evaluating the relevance of each document to the summary. Thus, a rationale for the decision can be made available, which marks the difference between LLM-based and human annotations. While traditional human annotation through crowdsourcing platforms like Amazon Mechanical Turk usually produces annotation results without underlying reasons due to additional costs, LLM-based annotators can easily offer explanations through CoT. These rationales can assist human managers in reviewing results and rectifying erroneous decisions.
@@
After annotation, we found that 27,052 of the 153,091 articles can be considered noisy documents and do not contribute to the summarization. Subsequently, we constructed Multi-News+ by removing these noisy documents from Multi-News while preserving the train/valid/test split. Figure 2 presents the comparison of the Multi-News and Multi-News+ datasets in terms of the number of documents per set. More than 15% of the documents in Multi-News are irrelevant, diminishing the dataset’s quality and degrading the model’s performance. Furthermore, 379 sets have no relevant source articles, as shown in Appendix G. In contrast, by deleting noisy documents, Multi-News+ demonstrates enhanced quality.
@@ 4.2 Result
The results in Table 2 demonstrate the superiority of the Multi-News+ dataset in enhancing the performance of summarization models compared to the original Multi-News dataset. Across various metrics, models trained on Multi-News+ consistently outperform those trained on Multi-News, indicating better summarization quality with the refined dataset. This highlights the effectiveness of dataset cleansing in removing noisy and irrelevant documents, thereby enhancing the overall performance of summarization models. Additionally, we performed a human evaluation on the output of 379 sets that are classified as having no relevant source articles and found that 356 sets are correctly classified. We provide an example of error analysis in Appendix H.
@@ 5 Conclusion
We acknowledge several limitations regarding our proposed method. First, our method is primarily limited by the possibility of wrong classification even with majority voting and CoT. In the future, we may adopt various LLMs as agents and apply weighted majority voting according to their performance to alleviate this issue.
@@
As we are exploiting LLMs for classifying irrelevant documents rather than text generation, the ethical concern with our method is smaller than that of studies that utilize LLMs to generate texts. Nonetheless, recent studies suggest that the CoT technique may induce ethical bias in LLM Shaikh et al. (2023). In future work, we plan to investigate this phenomenon’s appearance using our method.
@@ Appendix B Construction Process of Multi-News
However, this affected problems regarding the quality of the dataset. As shown in examples of noisy documents in Appendix F, several noisy documents consist of a message from Wayback Machine. Moreover, the failure to crawl the content of the webpage caused other problems. We investigated the case shown in Appendix G and found that it is a result of the crawling of the wrong part of the website. Figure 3 clearly showcases this phenomenon where the content in the red box is crawled instead of the content in the blue box, which is desired. Even though the content in the blue box is different for each article, the system wrongly crawled the shared red box, which resulted in five noisy documents that share the same content and do not contribute to the summary.
@@ Appendix C Implementation Details
We utilized PyTorch Paszke et al. (2019) and Huggingface Transformers Wolf et al. (2020) to implement and evaluate the model. Specifically, we employed facebook/bart-large-cnn444Note that this model is already fine-tuned with the CNN/DM dataset Nallapati et al. (2016), a single-document summarization dataset. and google-t5/t5-base, with 406M and 220M parameters, respectively, for BART and T5. Each model was trained using Adam Kingma and Ba (2015) with a learning rate of 2e-5 over 3 epochs. We used a batch size of 4 and implemented a gradient accumulation step of 4, resulting in a practical batch size of 16. For evaluation, we utilized bert-base-uncased and facebook/bart-large-cnn for BERTScore and BARTScore, respectively. We reported BERTScore-F1 in Table 2. ROUGE scores were measured using the rouge-score555https://pypi.org/project/rouge-score/ library, with the F1 score of each metric. The training was conducted on a single NVIDIA A100 40GB GPU. We provide the source code and dataset to the public.666https://anonymous.4open.science/r/MULTINEWS_PLUS/
@@
Appendix D Additional Experiment with Large Language Models
@@
Appendix E Analysis of Multi-News
@@
Appendix F Examples of Noisy Documents
@@
Appendix G Extreme Cases of Noisy Documents
@@
Appendix H Error Analysis

### v3

@@ Abstract
In this study, we present a case study that extends the application of LLM-based data annotation to enhance the quality of existing datasets through a cleansing strategy. Specifically, we leverage approaches such as chain-of-thought and majority voting to imitate human annotation and classify unrelated documents from the Multi-News dataset, which is widely used for the multi-document summarization task. Through our proposed cleansing method, we introduce an enhanced Multi-News+. By employing LLMs for data cleansing, we demonstrate an efficient and effective approach to improving dataset quality without relying on expensive human annotation efforts.
@@ 1 Introduction
The significance of dataset quality in deep learning applications cannot be overstated as mislabeled or noisy data can severely degrade performance Song et al. (2023). Datasets with incorrect labels, noise, or inconsistencies undermine the consistency and stability of model training. Cleansing these datasets contributes to enhancing model performance and generalization capabilities. Hence, ensuring the quality of the dataset by identifying and eliminating noisy data is imperative. In the realm of natural language processing, several researchers have attempted to improve the quality of noisy datasets Jiang et al. (2020, 2022). For example, ReDocRED Tan et al. (2022) addressed issues such as false negatives in DocRED Yao et al. (2019), a widely used dataset for relation extraction. Similarly, annotation inconsistencies were found in the MultiWOZ dataset Budzianowski et al. (2018) for dialogue state tracking Qian et al. (2021), leading to efforts to rectify these issues Eric et al. (2020); Zang et al. (2020); Han et al. (2021); Ye et al. (2022a).
@@ 3 Multi-News+
| Ablation Urlana et al. (2022) | 39.30 | 13.65 | 21.42 | 0.5967 | -2.457 |
The previous Multi-News dataset plays an important role in multi-document summarization research. It consists of sets of documents and their corresponding summaries. However, as shown in Table 1 and detailed in Appendix G and H, the Multi-News dataset contains several noisy and irrelevant articles that are unrelated to the summary or other documents. This issue arises from their construction process, which relies on automated crawling from the Internet Archive.
To solve this issue and cleanse the dataset, we defined our problem as a classification task determining whether each document is relevant to the summary. To this end, we designed the prompt for the model as shown in Appendix J. We integrated CoT to enhance the model’s performance by evaluating the relevance of each document to the summary. Thus, a rationale for the decision can be made available, which marks the difference between LLM-based and human annotations. While traditional human annotation through crowdsourcing platforms like Amazon Mechanical Turk usually produces annotation results without underlying reasons due to additional costs, LLM-based annotators can easily offer explanations through CoT. These rationales can assist human managers in reviewing results and rectifying erroneous decisions.
@@
After annotation, we found that 27,052 of the 153,091 articles can be considered noisy documents and do not contribute to the summarization. Subsequently, we constructed Multi-News+ by removing these noisy documents from Multi-News while preserving the train/valid/test split. Figure 2 presents the comparison of the Multi-News and Multi-News+ datasets in terms of the number of documents per set. More than 15% of the documents in Multi-News are irrelevant, diminishing the dataset’s quality and degrading the model’s performance. Furthermore, 379 sets have no relevant source articles, as shown in Appendix H. In contrast, by deleting noisy documents, Multi-News+ demonstrates enhanced quality.
@@ 4.2 Result
The results in Table 2 demonstrate the superiority of the Multi-News+ dataset in enhancing the performance of summarization models compared to the original Multi-News dataset. Across various metrics, models trained on Multi-News+ consistently outperform those trained on Multi-News, indicating better summarization quality with the refined dataset. This highlights the effectiveness of dataset cleansing in removing noisy and irrelevant documents, thereby enhancing the overall performance of summarization models. Additionally, we performed a human evaluation on the output of 379 sets that are classified as having no relevant source articles and found that 356 sets are correctly classified, which represents 93.9% of the human-machine agreement rate. We provide an example of error analysis in Appendix I.
Additionally, we conducted an ablation study using the cleansing method proposed by a previous study Urlana et al. (2022), detailed in Appendix F. Our findings indicate that this method is ineffective in improving downstream task performance on the Multi-News dataset, which focuses on multi-document summarization and differs from the configuration used in the prior study. This underscores the effectiveness of our proposed method and the value of Multi-News+.
5 Discussion and Future Works
In this section, we discuss recent advancements in the field since the submission of the manuscript and propose strategies for incorporating them in future research.
Cutting-edge models. Although we employed five GPT-3.5-turbo-0125 models for our experiments, the field has seen the release of more advanced models, such as GPT-4o OpenAI (2024b), GPT-4o-mini OpenAI (2024a), and OpenAI O1 OpenAI (2024c), along with the continued development of open-source models like LLaMA-3 Dubey et al. (2024), Gemma-2 Team et al. (2024), and Mistral Nemo Mistral (2024). Models such as GPT-4o-mini and other open-source alternatives offer reduced costs compared to GPT-3.5-turbo-0125, making their adoption promising for both lowering the expense of dataset cleansing and improving the accuracy of detecting noisy documents.
Weighted majority voting. The availability of high-performance yet cost-effective models like GPT-4o presents the opportunity to use them as expert annotators, given their superior capabilities compared to models like GPT-3.5-turbo-0125 or GPT-4o-mini. For example, rather than using five GPT-3.5-turbo-0125 models, we could employ three GPT-3.5-turbo-0125 models alongside one GPT-4o, with GPT-4o carrying double the weight of a GPT-3.5-turbo-0125 annotator. This approach positions GPT-4o as an expert, where agreement between at least one GPT-3.5-turbo-0125 model and GPT-4o would trigger document deletion.
Supervision from superior models. Another potential approach involves using more capable models to verify annotation results. In this scenario, GPT-4o would not participate in the initial annotation process but would instead verify the outcomes produced by GPT-3.5-turbo-0125 models. By taking the documents, summaries, and annotation results as input, GPT-4o acts as an expert reviewer overseeing the outputs of standard annotators.
Cost-efficient cleansing via pre-screening. In this paper, we applied the data cleansing strategy to every document in the dataset. However, a more cost-efficient approach could involve performing the annotation procedure only on documents likely to contain noise. Techniques such as dataset cartography Swayamdipta et al. (2020) could serve as a pre-screening method to identify cleansing candidates, thereby reducing the overall cost of dataset cleansing.
@@ 6 Conclusion
We acknowledge several limitations regarding our proposed method. First, our method is primarily limited by the possibility of wrong classification even with majority voting and CoT. In the future, we may adopt various LLMs as agents and apply weighted majority voting according to their performance to alleviate this issue, as discused in Section 5.
@@
As we are exploiting LLMs for classifying irrelevant documents rather than text generation, the ethical concern with our method is smaller than that of studies that utilize LLMs to generate texts. Nonetheless, recent studies suggest that the CoT technique may induce ethical bias in LLM Shaikh et al. (2023). In future work, we plan to investigate this phenomenon’s appearance in our method.
Acknowledgements
This research was supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(NRF-2022R1C1C1008534), and Institute for Information & communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program, Chung-Ang University).
@@ Appendix B Construction Process of Multi-News
However, this affected problems regarding the quality of the dataset. As shown in examples of noisy documents in Appendix G, several noisy documents consist of a message from Wayback Machine. Moreover, the failure to crawl the content of the webpage caused other problems. We investigated the case shown in Appendix H and found that it is a result of the crawling of the wrong part of the website. Figure 3 clearly showcases this phenomenon where the content in the red box is crawled instead of the content in the blue box, which is desired. Even though the content in the blue box is different for each article, the system wrongly crawled the shared red box, which resulted in five noisy documents that share the same content and do not contribute to the summary.
@@ Appendix C Implementation Details
We utilized PyTorch Paszke et al. (2019) and Huggingface Transformers Wolf et al. (2020) to implement and evaluate the model. Specifically, we employed facebook/bart-large-cnn444Note that this model is already fine-tuned with the CNN/DM dataset Nallapati et al. (2016), a single-document summarization dataset. and google-t5/t5-base, with 406M and 220M parameters, respectively, for BART and T5. Each model was trained using Adam Kingma and Ba (2015) with a learning rate of 2e-5 over 3 epochs. We used a batch size of 4 and implemented a gradient accumulation step of 4, resulting in a practical batch size of 16. For evaluation, we utilized bert-base-uncased and facebook/bart-large-cnn for BERTScore and BARTScore, respectively. We reported BERTScore-F1 in Table 2. ROUGE scores were measured using the rouge-score555https://pypi.org/project/rouge-score/ library, with the F1 score of each metric. The training was conducted on a single NVIDIA A100 40GB GPU. We provide the source code and dataset to the public.666https://github.com/c-juhwan/multi_news_plus
@@
Appendix D Manual Analysis
To perform a more detailed analysis of the accuracy of the proposed method, we randomly selected 60 instances from the validation set, which comprises 153 documents. A confusion matrix was defined to evaluate the classification for each document as follows:
True Positive (TP): Relevant documents that were correctly classified as relevant.
False Positive (FP): Documents classified as relevant but are not actually relevant.
True Negative (TN): Irrelevant documents correctly classified as not relevant.
False Negative (FN): Relevant documents incorrectly classified as not relevant.
Upon review, we found that 127 documents were classified as TP, 24 as TN, and 2 as FN. The annotation framework identified 26 documents as irrelevant and noisy, which accounts for approximately 17% of the total 153 documents. This aligns closely with the statistics in Table 3 of Appendix A, which indicates that 18.6% of documents in the validation set were classified as noisy.
From these results, the precision is 1.0, as there were no FP documents, while the recall is approximately 0.984. Additionally, we observed that 17 of the 24 TN documents could be classified as noisy system messages, such as “This will appear next to all of your comments; this will not appear anywhere on Newser,” as illustrated in Appendix G. The remaining 7 documents were irrelevant to the summary.
Furthermore, we investigated the two FN cases. In one instance, the summary included a portion related to the misclassified document at the very end. In the other, the misclassified document provided context for the summary but was not directly connected to it. These cases are consistent with the error patterns discussed in Appendix I.
It is important to note that while individual annotators occasionally made incorrect classifications, the majority voting process effectively corrected these errors. This highlights the efficacy of our proposed method in improving data annotation quality and ensuring thorough dataset cleansing.
Appendix E Additional Experiment with Large Language Models
@@
Appendix F Analysis of Multi-News
@@
Appendix G Examples of Noisy Documents
@@
Appendix H Extreme Cases of Noisy Documents
@@
Appendix I Error Analysis